# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v1

@@ 2.2 Data Reduction
Insight-HXMT, China’s first astronomical X-ray satellite, was launched on 2017 June 15 (Zhang et al. 2014, 2020) to observe and study the properties of black holes, neutron stars (NSs), active galactic nuclei (AGNs), and other astronomical phenomena. The main scientific instrument is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2. The satellite has three instruments or payloads: High Energy (HE; Liu et al. 2020), Medium Energy (ME; Chen et al. 2020), and Low Energy (LE; Cao et al. 2020). They have an effective area of 5100, 952, and 384 cm2 and cover an energy range of $20-250$, $5-30$, and $1-15$ keV, respectively.
After downloading on-demand level-1 data from the archive, we produced cleaned level-2 data for science analysis. This process of cleaning raw data was done in the following way. We first installed the HXMTDAS [^1] http://hxmt.org/index.php/usersp/dataan v2.05. Using this software, we run the hpipeline command using proper input and output directories. This pipeline performs several subsequent automatic commands for all three instruments. However, some suitable conditions are needed to be provided. Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8~{}GeV$; pointing offset angle $<0.04^{\circ}$; $>600~{}s$ away from the South Atlantic Anomaly (SAA). All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products. This is discussed in the HXMT manual [^2] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details. The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE. ME, and LE instruments. Whereas, the helcgen, melcgen, and lelcgen commands produce the light curve files for the three instruments. herspgen, merspgen, and lerspgen produce suitable response files. The background subtraction for both the timing and spectral data was performed using the commands hebkgmap, mebkgmap, and lebkgmap for instruments HE, ME, and LE, respectively. For $\chi^{2}$ fit-statistics in XSPEC, we group the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin. We also set the time bin size to $0.01s$ to produce suitable light curves for generating PDS and looking for QPOs.
Using the cleaned spectra and light curves, we analyze the source properties, which are discussed in detail in the very next subsection.
@@ 2.3 Data Analysis
We perform timing and spectral analysis of the very first outburst during 2023 of the BHC Swift J1727.8-1613. We have downloaded the daily average light curve using the MAXI/GSC [^3] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html public archive. Using light curves from all three modules of the HXMT, i.e., LE, ME, and HE, we have produced $0.01~{}s$ time-binned light curves first. Using those light curves, we have produced power density spectrum (PDS) by using the fast Fourier transformation (FFT) in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral will give the rms-squared fractional variability. White noise level is subtracted and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. Using these steps, we searched for LFQPOs. We fit the QPO shape by the use of a Lorentzian model in powspec and extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, normalization. For many observations, we have found the presence of single or multiple harmonics. We have also extracted their properties by fitting them with a Lorentzian model. We fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For spectral analysis, we have also used all three modules, i.e., LE, ME, and HE and fitted the broadband data in the $2-150$ keV energy band. Looking for the best fit, we have found that using the combination of disk blackbody, power-law, gaussian, and pexrav models, the data is best fitted. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we are simultaneously fitting all three modules, we have included a constant to normalize the three resultant fittings. The following two are our model combination: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was required), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for those for which gabs was not required. The spectral analysis was not done for all the exposures of the listed observations. Like the timing analysis, we did not spectrally analyze all the exposures, given in Table A. The spectrally analyzed exposures are marked with ‘*’ in Table A. The reason behind this is that, for timing analysis, we found variations of timing properties within a single day, whereas, for spectral analysis, the properties do not vary that much within such a short time. We report them in the result section.
We note that light curves for the LE instrument were not produced for some exposure IDs although we run the same pipeline for all three instruments. Thus we only represented those exposures MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
@@ 3 Results
Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties of the source using Insight-HXMT data. We broadly discuss the results in the two subsections below.
@@ 3.1 Temporal Properties
3.1.1 Outburst Profile and Hardness Ratio
In Figure 1, we show the variation of the flux during the outburst. The outburst started roughly at around MJD 60180 (2023 August 24), when its flux came out of quiescence. The MAXI/GSC flux started to rise after this date, as we can notice from panel (a) of Figure 1. Within five days, the flux increased very rapidly and the total flux reached its peak on MJD 60185 (2023 August 29). Then the flux started to decrease very slowly. It decreased very slowly and gradually except for the period from MJD 60200 to 60225, where the flux showed a significant variation. In panel (b), we show the variations of the $2-6$ and $6-20$ keV energy bands count rates with time. We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly and gradually without showing any other variation, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux. MJD 60185 is the peak for both the total $2-20$ keV and hard $6-20$ keV fluxes, whereas, the soft flux reached its peak much later at around MJD 60217. This suggests that the variation in the total count rate is mainly due to the variation of the soft component. After this period, the soft flux also decreased slowly to reach the end of the outburst and enter again into the quiescence phase.
In the panel (c), we show the variation of the hardness ratio, which is the ratio of the hard $6-20$ keV to the soft $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$. From this value, it first gradually decreased to MJD 60235 as the outburst progressed. After that, it became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of our taken data of the outburst on MJD 60350. The hardness ratio is defined as $F_{H}/F_{S}$, where $F_{H}$ and $F_{S}$ are the hard and soft fluxes respectively. When this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick and rough idea about the spectral nature of an outburst. However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux. However, after a few days, soft flux completely took over the hard flux, and the HR was also not high. Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst. After MJD 60185, it made a transition into the SS. Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification. We only spectrally define the source in the duration which falls into our analysis range.
@@ 3.1.2 Quasi Periodic Oscillations (QPOs)
As described in §2, we used the $0.01s$ time-binned light curves from all three bands (LE, ME, and HE) for our timing analysis. Using those light curves, we produced the power density spectrum (PDS) of those light curves to study QPOs. We find that QPO nature was present in all of the light curves in all three energy bands. In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs that we have used for our timing analysis. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 2). While we fit every exposure ID of this obs ID, we found that the QPO frequency was changing even within this small period. This is why we fitted all the light curve data of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
@@
Refer to captionRefer to caption
@@
In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$. Then it decreased for a very short period until it very rapidly increased to $1.5$Hz again in all three bands. After that, it showed varying nature till MJD $\sim 60197$ and then it finally increased again. This might be caused by the varying nature of the origin of these oscillations. We will discuss this in the next section.
Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs. We also extracted source and background count rates from the exposures. Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. The values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
@@
In Figure 4, we show a 3-D representation of the evolution of QPO RMS with corresponding frequency. In the LE band (Fig.4a), The RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 4b), it increased from the onset of the analysis period for some time and then again decreased. HE band (Fig. 4c) shows a similar trend as of LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we would not put much insight into the classification of the QPO nature. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposure IDs that have been presented here.
@@ 3.2 Spectral Properties
Along with studying the timing properties, studying the spectral properties gives a clearer idea about the nature of the source. With the available Insight-HXMT data, we analyzed the source for a total of 32 exposure IDs. These exposure IDs are marked with a ‘*’ sign in the 1st column of Table 2. This source hasn’t been extensively analyzed spectrally. We perform an extensive study of this source’s spectral properties for the first few days after the onset of its outburst. We started our spectral analysis on MJD 60181.4. We have used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all these exposure IDs.
@@
Figure 5: Model fitted spectrally analyzed unfolded spectra. (a) is or observation ID. P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was necessary to achieve the best fit. (b) is for the observation ID. P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not needed to achieve the best fit.
We first fitted the data with phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposure, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 5b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
Model Fitted Insight-HXMT Spectra
@@ 3.3 Correlation between spectral and timing properties
The QPO frequency could be produced from the oscillation of the Compton corona, which could also be the origin of the hard power-law photons that extend the spectrum to very high energies. They could be correlated to each other. It has been previously seen that the QPO frequency ($\nu_{qpo}$) and the photon index of power-law ($\Gamma$) show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 7, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as below.

### v2

@@ 2.2 Data Reduction
The main scientific instrument of Insight-HXMT is an array of 18 NaI/CSI phoswich scintillation detectors, each of which covers an effective area of 286 cm2. The satellite has three instruments or payloads: High Energy (HE; Liu et al. 2020), Medium Energy (ME; Chen et al. 2020), and Low Energy (LE; Cao et al. 2020). They have an effective area of 5100, 952, and 384 cm2 and cover an energy range of $20-250$, $5-30$, and $1-15$ keV, respectively.
After downloading on-demand level-1 data from the archive, we produced cleaned level-2 data for science analysis. The processing of cleaning raw data was done in the following way. We first installed the HXMTDAS [^2] http://hxmt.org/index.php/usersp/dataan v2.05. Using this software, we run the hpipeline command using proper input and output directories. This pipeline performs several subsequent automatic commands for all three instruments under some given conditions. Some conditions were set to get a good time interval, e.g., elevation angle $>10^{\circ}$; geomagnetic cutoff rigidity $>8$ GeV; pointing offset angle $<0.04^{\circ}$; $>600$ sec away from the South Atlantic Anomaly (SAA). All these commands altogether extract and clean the raw data and produce cleaned and analyzable science products [^4] This is discussed in the HXMT manual [^3] http://hxmten.ihep.ac.cn/SoftDoc/501.jhtml in full details. The specific commands hespecgen, mespecgen, and lespecgen produce the spectra for HE, ME, and LE instruments. Whereas, the helcgen, melcgen, and lelcgen tasks produce the light curve files for the three instruments. herspgen, merspgen, and lerspgen produce corresponding response files. The background subtraction for both the timing and spectral data was performed using the commands hebkgmap, mebkgmap, and lebkgmap for instruments HE, ME, and LE, respectively. For the $\chi^{2}$ fit-statistics in XSPEC, we grouped the spectrum using the grppha task of FTOOLS to a minimum of 30 counts per bin. We also set the time bin size to $0.01$ sec to produce the light curves for generating PDS and QPOs.
Using the cleaned spectra and light curves, we further fitted and analyzed to extract the outburst properties, which are discussed in the next section.
@@ 2.3 Data Analysis
We have downloaded the daily average light curve of MAXI/GSC [^5] http://maxi.riken.jp/star_data/J1727-162/J1727-162.html from the archive to study the outburst profile in addition to Insight-HXMT. Using the unbinned light curves from LE, ME, and HE modules, we have first produced $0.01$ sec time-binned light curves. Then the power density spectrum (PDS) was generated using the fast Fourier transformation (FFT) technique in the powspec task of the XRONOS package in the HEASoft software. Each observation’s data was subdivided into several intervals, each of which contains 8192 newbins. First, the PDS for each interval is generated and then they are averaged to make a resultant PDS. The PDS is normalized in a way that their integral can give the RMS-squared fractional variability. The white noise level is subtracted and a geometrical rebinning of -1.02 is used. For some data, we needed to use a geometrical rebinning of -1.05 to have a prominent QPO-like nature in the PDS. We fit the QPO using single or multiple Lorentzian model components in powspec to extract QPO properties like frequency ($\nu_{qpo}$), full-width at half maximum or FWHM, and normalization. For several observations, we have found the presence of single or multiple harmonics. We have also extracted their properties using the same fitting method. We have fitted the light curves of all the exposures of the listed 10 observations. We report them in the result section.
For the spectral analysis, we have used the same LE, ME, and HE modules and fitted the broadband data in the $2-150$ keV energy band. The best is achieved by using the combination of disk blackbody, power-law, gaussian, and pexrav models. For some observations, we needed to add a gabs model to achieve the best fit. For interstellar absorption, we have used the tbabs model. Since we simultaneously fit all three modules, we have included a constant to normalize the three resultant fittings. The following two model combinations are used: i) constant*tbabs*gabs*(diskbb + power-law + gaussian + pexrav), ii) constant*tbabs*(diskbb + power-law + gaussian + pexrav) (for which gabs was not required. The spectral analysis was done for the exposures are marked with ‘*’ in Table 2. We did not do the spectral analysis for all epochs as spectral variability was not present within such a small timescale (intraday) like the variability in the light curves. We report them in the result section.
We note that light curves for the LE instrument were not produced for some exposures. Therefore, we represented only those exposures’ MJDs in Table 3 for which all three light curves were available for analysis for uniformity.
@@ 3 Results
Using the timing and spectral analysis, we have studied the accretion flow properties of the very recent 2023 outburst of the BHC Swift J1727.8-1613. We studied the QPO properties of the source during the outburst as well as the spectral nature and radiation properties using Insight-HXMT data. We broadly discuss the results in the next two sections.
@@ 3.1.1 Outburst Profile and Hardness Ratio
Figure 1 shows the variation of the flux obtained from MAXI/GSC during the outburst. The outburst started roughly at around MJD 60180 (2023 August 24), when its flux came out of quiescence. The MAXI/GSC flux started to rise after this date, as we can notice from panel (a) of Figure 1. Within five days, the flux increased very rapidly and the total flux reached its peak on MJD 60185 (2023 August 29). After that, the flux started to decrease very slowly, except for the period from MJD 60200 to 60225, where the flux showed a significant change. In panel (b), we show the variations of the count rates with time in the $2-6$ and $6-20$ keV energy bands. We notice that, after MJD 60185, the hard flux (i.e., $6-20$ keV) decreased very slowly, whereas, the soft $2-6$ keV flux showed a lot of variation just like the total $2-20$ keV flux. On MJD 60185 the $2-20$ keV and hard $6-20$ keV fluxes reached their peak values, whereas, the soft flux reached its peak much later at around MJD 60217. This suggests that the variation in the total count rate is mainly due to the variation of the soft component. After this period, the soft flux also decreased slowly till the end of the outburst and entered again into the quiescence phase.
In panel (c), we show the variation of the hardness ratio, which is the ratio of the hard ($F_{H}$) $6-20$ keV to the soft ($F_{S}$) $2-6$ keV flux. At the start of the outburst, the HR was $\sim 0.5$ and then it gradually decreased to MJD 60235 as the outburst progressed. After that, the HR became almost constant, although decreasing very slowly to $\sim 0.025$ at the end of the considered data of the outburst on MJD 60350. The way HR is defined, when this value is on the higher side, the spectral nature of the source is hard. This is the general case when an outburst is starting. When this value is low, there is dominance of soft flux over hard flux and the spectral nature would be soft. Without studying the detailed spectral analysis, HR provides a quick idea about the spectral nature of an outburst. However, for this source, we notice that HR was a little higher at the onset of the outburst, where the hard flux was also comparable to the soft flux. However, after a few days, soft flux completely took over the hard flux, and the HR was also not high. Depending on the HR value, we propose that the source was already in the SIMS after the onset of the outburst. After MJD 60185, it made a transition into the SS. Since we are not analyzing the full duration of the outburst, it is not the scope of the paper to study the full spectral classification. We only define the spectral state of the source based on the duration of which data is used in this work.
@@ 3.1.2 Quasi Periodic Oscillations (QPOs)
As described in §2, we used the $0.01$ sec time-binned light curves from all three bands (LE, ME, and HE) and produced the PDS to study QPOs. We find that QPO nature was present in all the light curves in all three energy bands. In Figure 2, we show the variation of the QPO frequency ($\nu_{qpo}$) for the 10 different obs. IDs as stated before. We first checked the first obs. ID P0614338001. It has several exposure IDs (given in Table 2). While we fit every exposure of this obs ID, we found that the QPO frequency was changing even within this small period. This is the reason we fitted all the light curves of the exposures of the obs IDs that we have taken for analysis. We find that for every single obs ID, the $\nu_{qpo}$ showed significant variation. Figure 2(a), (b), (c) show the variations of the $\nu_{qpo}$ with time for 10 different obs IDs for LE, ME, and HE respectively.
@@
Refer to captionRefer to captionRefer to caption
@@
In Figure 3, we show the evolution of the $\nu_{qpo}$ for the total duration of our analysis period. In (a), (b), and (c) panels the variations are for LE, ME, and HE respectively. We find that the QPO frequencies were $\sim 0.23\pm 0.008$, $0.24\pm 0.002$, and $0.24\pm 0.002$ Hz for LE, ME, and HE on the starting day of our analysis. After this day, the $\nu_{qpo}$ shows a rapid increase with time till $\sim$ MJD 60187 when $\nu_{qpo}$ was $\sim 1.2$Hz for all the three bands. After this date, it increased very slowly till MJD $\sim 60190$ when $\nu_{qpo}$ was $\sim 1.45$. Then it decreased for a very short period until it rapidly increased to $1.5$Hz again in all three bands. After that, it showed varying nature till MJD $\sim 60197$, and then it finally increased again. We have discussed the possible reason behind this nature in the next section.
Using the PDS fitting, we extracted some information (e.g., full-width at half maximum (FWHM), Normalization (LN)) of QPOs. We also extracted source and background count rates for the exposures. Using these estimates, we calculated the $Q$-value and RMS that represent the sharpness of the QPO and the fractional variability in the PDS, respectively. These values are listed in Table 3 in columns 5–7 ($Q$-value) and 8–10 (RMS) for LE, ME, and HE respectively.
@@
In Figure 4, we show a 3-D representation of the evolution of QPO and RMS with time. In the LE band (Fig. 4a), the RMS was higher at the start of the analysis period and it then gradually decreased. In the ME band (Fig. 4b), it increased from the onset of the analysis period for some time and then again decreased. The HE band (Fig. 4c) shows a similar trend as in LE band. Since the QPO nature has already been reported in the recent paper by Yu et al. (2024), we are not focusing on the classification of the QPO nature. However, by the values of both $Q$-factor and RMS, it can be said that the QPO is type-C in nature for all the exposures reported here.
@@ 3.2 Spectral Properties
Along with the timing properties, studying the spectral properties provides more insights on the nature of the outburst. With the available Insight-HXMT data, we analyzed the source for a total of 32 exposures. These exposure IDs are marked with a ‘*’ sign in the 1st column of Table 2. An extensive spectral analysis on this source has not been done so far. We perform a detailed spectral analysis on this source for the first time using HXMT data. We started our spectral analysis from MJD 60181.4. We have simultaneously used LE+ME+HE in the $2-150$ keV energy band (LE in 2–10, ME in 10–35, and HE in 35–150 keV) for our spectral fitting for all the selected exposure IDs.
@@
Figure 5: Model fitted spectrally analyzed unfolded spectra. (a) is or observation ID. P061433800501 (Exposure: P061433800501-20230902-01-01) for which the gabs model was required to achieve the best fit. (b) is for the observation ID. P061433800601 (exposure: P061433800601-20230904-01-01) for which the gabs model was not required to achieve the best fit.
We first fitted the data with phenomenological diskbb and power-law models. We used the multiplicative tbabs model to account for the interstellar absorption. Our model read as constant*tbabs(diskbb + powerlaw) in XSPEC. The constant is taken to normalize between three different instruments in LE, ME, and HE. We found that the fit was not statistically acceptable. There was a clear sign of the Gaussian feature in the unfolded spectrum around $5-6$ keV. Also, there was a hump-like feature $\sim 20$ keV. Besides these features, the absorption feature was noticed below 2 keV. Thus, we refitted the data using the model combination as constant*tbabs*gabs(diskbb + power-law + ga + pexrav). Here, the gabs model takes care of the low energy absorption feature, while the ga and pexrav models take care of the contribution from the Gaussian and hump-like nature, as mentioned above. In Fig. 5a, we show the best-fitted unfolded spectrum using this model combination for the observation ID P061433800501 (Exposure ID: P061433800501-20230902-01-01), where $\chi^{2}/DOF\sim 1$. After some exposure, we found that the gabs model was not needed anymore to achieve the best fit. Then, the model combination in XSPEC reads as constant*tbabs(diskbb + power-law + ga + pexrav). In Figure 5b, we show an unfolded best-fitted spectrum using the later model combination. This is for the observation ID. P061433800601 (exposure ID: P061433800601-20230904-01-01), for which $\chi^{2}/DOF\sim 0.9$.
Model Fitted Parameters
@@ 3.3 Correlation between spectral and timing properties
The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies. Therefore, these two features might be interlinked or correlated. It has been previously seen that the QPO frequency ($\nu_{qpo}$) and $\Gamma$ show a positive correlation for some of the sources during their active outbursting phase (Vignarca et al. 2003; Shaposhnikov & Titarchuk 2009; Stiele et al. 2013). We also tried to find out if there are any correlations between these two properties. In Figure 7, we show the variations of the photon index with the QPO frequency for the three energy bands. Considering these points, we have found a positive correlation between $\Gamma$ and $\nu_{qpo}$. Using both the Pearson-Linear and Spearman-Rank correlation methods, we found correlation coefficients for the three energy bands as shown in the table below.